You have prepared 250mL of a 0.200M acetate buffer solution with a pH of 4.53.

b. If you made the solution above using solid sodium acetate (MW 136 g/mol) and liquid acetic acid (17.6 M, referred to as glacial acetic acid), what mass of sodium acetate is required and what volume of glacial acetic acid is required? Show all your calculations. mass of sodium acetate volume of glacial acetic acid
Answer #1

total moles of buffer = 250 x 0.2 / 1000 = 0.05

acetic acid (x) + sodium acetate (y) = 0.05 ---------------------------> (1)

pH = pKa + log [y / x]

4.53 = 4.74 + log (y / x)

y/ x= 0.617

y = 0.617 x-----------------------> (2)

by solving 1 and 2

x = 0.0309

y = 0.0191

mass of sodium acetate = moles x molar mass

                                    = 0.0191 x 136

                                    = 2.6 g

volume of acetic acid = moles / molarity

                                = 0.0309 / 17.6

                                 = 4.07 x 10^-3 L

                                = 4.07 mL